1 litre of glass bulb contains 2*10^21  molecules at a pressure of   7.57*10^3  N/m^2     find the RMS and the most probable velocities  and temparature of the gas //???????
 

Grade:10

1 Answers

Vikas TU
7 years ago
Vrms = root(3RT/M) ….............(1)
Now from ideeal gas eqn. we get,
7.57*10^3  * 1 = 2*10^21/6*10^23 RT.................(2)
From eqn. (2) and (1) we get,
with M = 28 gm (Nitrogen gas molar mass).
Vrms = 493.276 m/s
and 
T = 276.61 x 10^5 x 10^-5 => 276.61 K.